OpenStudy (anonymous):

(x+3)(2x-5) simplify

OpenStudy (anonymous):

if you could tell me how to do these it would be much appriciated >-<

OpenStudy (anonymous):

(a+b)(c+d)=ac+ad+bc+bd So, x*2x-5x+6x-15=\[2x ^{2}+x-15\] that's what i get from 'simplify'